Using the ε −δ definition, show that lim 3( x )5 1
1
Expert's answer
2020-05-12T17:28:22-0400

We have to prove that given any "\u03b5>0"  we can find a "\u03b4_\u03b5>0"  such that

"|
f
(
x
)
−
1
|
<
ε
 \quad \forall
x
∈
(
2
−
δ_
ε
,
2
+
δ_
ε
)"

Let's evaluate the difference

"|\nf\n(\nx\n)\n\u2212\n1\n|\n=\n|\n3\nx\n\u2212\n5\n\u2212\n1\n|\n=\n|\n3\nx\n\u2212\n6\n|\n=\n3\n|\nx\n\u2212\n2\n|"

Therefore

"|
f
(
x
)
−
1
|
<
ε
⇔
|
x
−
2
|
<
\frac{ε}{3}"

Given, that "\u03b5>0" we can choose "δ_ε<\frac{ε}{3}"

Hence

"x
∈
(
2
−
δ_
ε
,
2
+
δ_
ε
)
⇒
|
x
−
2
|
<
\frac{ε}{3}
⇔
|
f
(
x
)
−
1
|
<
ε"